# Exploiting Agent Symmetries for Performance Analysis of Distributed Optimization Methods

### v1

@@ 4.1 The PEP formulation restricted to fully symmetric solutions
When all agents are equivalent (Assumption 1), Corollary 6.1 shows that it does not impact the value of the worst-case but we will see in Theorem 7 that it allows writing the PEP in a compact form since the performance measure and the constraints can be expressed only in terms of the smaller blocks $f_{A}$, $G_{A}$, and $G_{B}$ that are repeating in $F^{s}$ and $G^{s}$. Theorem 7 states sufficient conditions for which a PEP for distributed optimization can be made compact and for which it provides a worst-case value independent of the number of agents $n$. The theorem relies on the following definitions.
@@
The variables common to all agents, such as $\overline{x}$ or $x^{*}$, can indeed be used in scale-invariant expression because they are assumed to be copied in each agent set of variables:
@@
[…]
@@
[…]
@@
As detailed in Section 2.2, the set of averaging matrices $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ has Gram-representable necessary interpolation constraints, see Theorem 4 applied to the consensus subspace $\mathcal{C}$. Since $\mathcal{A},\mathcal{P},\mathcal{I}$ and interpolation constraints for $\mathcal{F}$ are also Gram-representable, Proposition 2 guarantees that the computation of $w(n,\mathcal{A},t,\mathcal{P},\mathcal{I},\mathcal{F},\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}})$ can be formulated as an agent-dependent SDP PEP problem (14). Moreover, since the agents are equivalent, we can restrict to fully symmetric solutions without impacting the worst-case value (see Corollary 6.1). Hence, all the PEP components can be written in terms of $F^{s}$ and $G^{s}$, which are composed of blocks $f_{A}$, $G_{A}$, and $G_{B}$. Lemma 7.1 shows how the SDP constraint $G^{s}\succeq 0$, can be expressed in terms of $G_{A}$ and $G_{B}$. The other elements of the PEP can directly be expressed in terms of $f_{A}$, $G_{A}$, and $G_{B}$, using the definition of $F^{s}$ and $G^{s}$ (77) (81), as detailed in Appendix A. Since all the PEP elements can be expressed in terms of blocks $f_{A}\in\mathbb{R}^{q}$, $G_{A}\in\mathbb{R}^{p\times p}$ and $G_{B}\in\mathbb{R}^{p\times p}$, the problem can be made smaller by considering them as the variables of the SDP PEP, instead of the full matrix $G^{s}\in\mathbb{R}^{np\times np}$ and vector $F^{s}\in\mathbb{R}^{nq}$. The size of the problem is thus independent of the number of agents $n$, but $n$ could still appear as a coefficient in the constraints or the objective of the problem. ∎
@@
[…]
@@
The new variable $G_{C}$ allows writing many constraints without any dependence on $n$. Let us state a lemma, that will be used to prove the second part of Theorem 7.
@@
By Corollary 6.1, since all the agents are equivalent, we can restrict the PEP to fully symmetric solution $(F^{s},G^{s})$ from (77)-(81). Since the expression is scale-invariant and Gram-representable (Definition 10), it can only combine three types of terms, with coefficients independent of $n$,
@@
    where $e_{f(x)}$ is the vector of coefficients such that $e_{f(x)}^{T}f_{i}=f_{i}(x_{i})$. The second equality holds because $f_{i}=f_{A}$ for all $i$, by definition of the agent-symmetric $F^{s}$ (77).
@@
    The average over the $n^{2}$ pairs of agents of the scalar products between two variables each related to any agent:
@@
Then, we prove that, in the symmetrized PEP, any single-agent constraint (Definition 11) can be written with $f_{A}$ and $G_{A}$, independently of $n$. Since the PEP components are all Gram-representable, any of its single-agent constraints for agent $i$ linearly combines local function values $f_{i}(x_{i})$ and scalar products of local variables $x_{i}^{T}y_{i}$, with coefficients that are independent of $n$. Here, $x_{i}$ and $y_{i}$ denote any local vector held by agent $i$. By definition of the symmetric $F^{s}$ (77), a local function value term can be expressed with $f_{A}$ as
@@
which is well scale-invariant thanks to the $1/n$ factor. Finally, the interpolation constraints for consensus steps with averaging matrices from $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ are given by (50), (51) and (52), from Theorem 4 applied to the consensus subspace $\mathcal{C}$. In these constraints, notations hide sums over the agents and can be written with Gram-representable and scale-invariant expressions as
@@
In the end, all the PEP components can well be expressed independently of $n$, with $f_{A}\in\mathbb{R}^{q}$, $G_{A}\in\mathbb{R}^{p\times p}$ and $G_{C}\in\mathbb{R}^{p\times p}$. We refer to Appendix A for explicit expressions of different possible PEP constraints and objectives. Therefore, the SDP PEP can be made smaller and fully independent of the number of agents $n$, by considering these blocks as variables of the problem, instead of the full matrix $G^{s}\in\mathbb{R}^{np\times np}$ and vector $F^{s}\in\mathbb{R}^{nq}$. ∎
@@ 4.2 Impact of agent symmetry on the worst-case solution
    $x_{i}^{k}=R_{i}x_{1}^{k}\qquad\text{with $R_{i}\in\mathbb{R}^{d\times d}$ suchthat $R_{i}^{T}R_{i}=I$,}\hskip 56.9055pt\text{for all $i\in\mathcal{V}$ and all $k$}.$
@@
$\quantity[P_{i}\leavevmode\nobreak\ P_{j}]=\tilde{R}_{ij}\quantity[P_{1}\leavevmode\nobreak\ P_{2}]\qquad\text{with $\tilde{R}_{ij}\in\mathbb{R}^{d\times d}$ such that $\tilde{R}_{ij}^{T}\tilde{R}_{ij}=I$,}\hskip 42.67912pt\text{for all $i,j\in\mathcal{V}$ and all $k$}.$
@@
By Corollary 6.1, when the agents are all equivalent, there is a worst-case solution of the agent-dependent PEP, which is fully symmetric: $(F,G)=(F^{s},G^{s})$. When the solution is agent-symmetric, by definition of $G^{s}$ (81), we have
@@
Concerning the local functions, by definition of $\mathcal{F}^{s}$ (77), we have, for all $i,j\in\mathcal{V}$,
$f_{i}=f_{j}\quad\text{where $f_{i}\in\mathbb{R}^{q}$ is a vector containing the functions values related to agent $i$.}$ (103)
@@ 5 Compact PEP with multiple equivalence classes of agents
While the agents are all equivalent (Assumption 1) for many common performance evaluation settings, there are advanced settings for which there are several equivalence classes of agents, see Definition 9. For example, when different groups of agents use different (uncoordinated) step-sizes, function classes, initial conditions, or even different algorithms. It can also happen that the performance measure focuses on a specific group of agents, e.g. the performance of the worst agent. The ability to efficiently and accurately evaluate the performance of distributed optimization algorithms in such advanced settings would enable a more comprehensive analysis and deeper understanding of algorithms performance.
@@
In this section, we exploit symmetries in different equivalence classes of agents in PEP, revealed in Proposition 6, to write agent-dependent PEPs in a compact form whose size only depends on the number of equivalence classes $m$, and not on the total number of agents $n$. By Proposition 6, we can solve the agent-dependent PEP problem, described in Section 2, restricted to symmetric agent-class solutions $(F^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$, without impacting the worst-case value. These symmetric solutions $(F^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$, defined in (72) and (76), depend on the partition $\mathcal{T}$ of the set of agents $\mathcal{V}$ into equivalence classes: $\mathcal{T}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{m}\}$, see Definition 9. A symmetrized solution $(F^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ has equal blocks for equivalent agents. In this section, we order the agents by equivalence classes in the symmetrized solution $(F^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ (72)-(76) so that we can write it as
@@
$F^{s}_{\mathcal{T}}$ $=\begin{bmatrix}F_{\mathcal{V}_{1}}&\dots&F_{\mathcal{V}_{m}}\end{bmatrix}^{T}$ where $\text{where }F_{\mathcal{V}_{u}}=\mathbf{1}_{n_{u}}\otimes f_{A}^{u}$ (105)
@@
$G^{s}_{\mathcal{T}}$ $=\begin{bmatrix}G_{\mathcal{V}_{1}}^{s}&G_{\mathcal{V}_{1}\mathcal{V}_{2}}^{s}&\dots&G_{\mathcal{V}_{1}\mathcal{V}_{m}}^{s}\\ G_{\mathcal{V}_{2}\mathcal{V}_{1}}^{s}&G_{\mathcal{V}_{2}}^{s}&&\vdots\\ \vdots&&\ddots&\\ G_{\mathcal{V}_{m}\mathcal{V}_{1}}^{s}&\dots&&G_{\mathcal{V}_{m}}^{s}\end{bmatrix}$ (110)
@@
and where $f_{A}^{u}\in\mathbb{R}^{q}$, $G_{A}^{u},G_{B}^{u},G_{E}^{uv}\in\mathbb{R}^{p\times p}$ for $u,v=1,\dots,m$ are the blocks repeating in $F^{s}_{\mathcal{T}}$ and $G^{s}_{\mathcal{T}}$, which have been defined in Section 3.
Remark (Limit cases for $m$).
@@
  • $m=n$: When no agents are equivalent, the partition $\mathcal{T}$ contains $m=n$ equivalence classes of size 1 and then the symmetrized agent-class solution $F^{s}_{\mathcal{T}}$, $G^{s}_{\mathcal{T}}$, defined in Proposition 6 is identical to the agent-dependent PEP solution $F$, $G$:
    $F^{s}_{\mathcal{T}}=F,\qquad\text{ and }\qquad G^{s}_{\mathcal{T}}=G.$
@@
    Therefore, the resulting PEP formulation is equivalent to the agent-dependent PEP from Section 2 and is not compact as its size is scaling with $m=n$.
@@
  • $m=1$: When all the agents are equivalent, the partition $\mathcal{T}$ contains $m=1$ equivalence class of size $n$ and then the symmetrized agent-class solution $F^{s}_{\mathcal{T}}$, $G^{s}_{\mathcal{T}}$ defined in Proposition 6 is identical to the fully symmetrized solution $F^{s}$, $G^{s}$ defined in Corollary 6.1
    $F^{s}_{\mathcal{T}}=F^{s},\qquad\text{ and }\qquad G^{s}_{\mathcal{T}}=G^{s}.$

### v2

@@ 4.1 The PEP formulation restricted to fully symmetric solutions
When all agents are equivalent (Assumption 1), Corollary 6.1 shows that restricting PEP to fully symmetric solutions does not impact the value of the worst-case. We will see in Theorem 7 that it allows writing the PEP in a compact form since the performance measure and the constraints can be expressed only in terms of the smaller blocks $f_{A}$, $G_{A}$, and $G_{B}$ that are repeating in $\mathbf{f}^{s}$ and $G^{s}$. Theorem 7 states sufficient conditions for which a PEP for distributed optimization can be made compact (i.e. with size independent of $n$) and for which it provides a worst-case value independent of the number of agents $n$.
Before stating the theorem, we introduce two types of expressions or constraints that we can allow in PEP to obtain agent-independent worst-case values. Let us consider a simple example to motivate the first type of expressions.
Example.
If we consider a PEP for distributed optimization with the initial conditions $f_{i}(x_{i}^{0})-f_{i}(x^{*})\leq 1$ for all $i\in\mathcal{V}$ and the performance measure $\mathcal{P}=\sum_{i=1}^{n}(f_{i}(x_{i}^{0})-f_{i}(x^{*})),$ then the worst-case value would scale with the number of agents $n$ in the problem : if we consider more agents, the maximum value of $\mathcal{P}$ will increase. To obtain a PEP that provides worst-case values independent of $n$, the performance measure needs to be an expression that is invariant to the number of agents, $\frac{1}{n}\sum_{i=1}^{n}(f_{i}(x_{i}^{0})-f_{i}(x^{*}))$ in our example. We propose a notion called scale-invariant, defined below.
@@
The variables common to all agents, such as $\overline{x}$ or $x^{*}$, can indeed be used in scale-invariant expressions because they are assumed to be copied in each agent set of variables :
@@
Let us consider another simple example to motivate the second type of expressions we define.
Example.
If we consider a PEP for distributed optimization with the initial conditions $\|x_{i}^{0}-x^{*}\|^{2}\leq n\quad\text{for all $i\in\mathcal{V}$}$ and the performance measure $\mathcal{P}=\frac{1}{n}\sum_{i=1}^{n}\|x_{i}^{0}-x^{*}\|^{2}$, then the worst-case value would scale with the number of agents $n$ in the problem : if we consider more agents, the maximum value of $\mathcal{P}$ will increase. To obtain a PEP that provides worst-case values independent of the number of agents $n$, the initial constraints involving only one agent at a time need to be independent of $n$, $\|x_{i}^{0}-x^{*}\|^{2}\leq 1$ in this example. We propose the notion of single-agent constraint, defined below.
@@
Let $n\geq 2$ and $\mathcal{W}=\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ with $\lambda^{-}\leq\lambda^{+}\in(-1,1)$. We assume that the algorithm $\mathcal{A}\in\mathcal{A}_{D}$, the performance measure $\mathcal{P}$, the initial conditions $\mathcal{I}$, and the interpolation constraints for $\mathcal{F}$ are linearly (or LMI) Gram-representable. If Assumption 1 holds, meaning that all the agents are equivalent in the PEP, then
@@
$G^{s}\succeq 0\qquad\Leftrightarrow\qquad G_{A}+(n-1)G_{B}\succeq 0\quad\text{and}\quad G_{A}-G_{B}\succeq 0.$ (83)
@@
As detailed in Section 2.2, the set of averaging matrices $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ has Gram-representable necessary interpolation constraints, see Corollary 4.1. Since $\mathcal{A},\mathcal{P},\mathcal{I}$ and interpolation constraints for $\mathcal{F}$ are also Gram-representable, Proposition 2 guarantees that the computation of $w(n,\mathcal{A},t,\mathcal{P},\mathcal{I},\mathcal{F},\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}})$ can be formulated as an agent-dependent SDP PEP problem (15). Moreover, since the agents are equivalent, we can restrict to fully symmetric solutions without impacting the worst-case value (see Corollary 6.1). Hence, all the PEP components can be written in terms of $\mathbf{f}^{s}$ and $G^{s}$, which are composed of blocks $f_{A}$, $G_{A}$, and $G_{B}$. Lemma 7.1 shows how the SDP constraint $G^{s}\succeq 0$, can be expressed in terms of $G_{A}$ and $G_{B}$. The other elements of the PEP can directly be expressed in terms of $f_{A}$, $G_{A}$, and $G_{B}$, using the definition of $\mathbf{f}^{s}$ and $G^{s}$ (76) (77), as detailed in Appendix A. Since all the PEP elements can be expressed in terms of blocks $f_{A}\in\mathbb{R}^{q}$, $G_{A}\in\mathbb{R}^{p\times p}$ and $G_{B}\in\mathbb{R}^{p\times p}$, the problem can be made smaller by considering them as the variables of the SDP PEP, instead of the full matrix $G^{s}\in\mathbb{R}^{np\times np}$ and vector $\mathbf{f}^{s}\in\mathbb{R}^{nq}$. The size of the problem is thus independent of the number of agents $n$, but $n$ could still appear as a coefficient in the constraints or the objective of the problem. ∎
@@
$G^{s}\succeq 0\qquad\Leftrightarrow\qquad G_{C}\succeq 0\quad\text{and}\quad G_{A}-G_{C}\succeq 0.$ (87)
@@
The new variable $G_{C}$ allows writing many constraints without any dependence on $n$. Let us state a lemma that will be used to prove the second part of Theorem 7.
@@
By Corollary 6.1, since all the agents are equivalent, we can restrict the PEP to a fully symmetric solution $(\mathbf{f}^{s},G^{s})$ from (76)-(77). Since the expression is scale-invariant and Gram-representable (Definition 11), it can only combine three types of terms, with coefficients independent of $n$,
@@
    where $e_{f(x)}$ is the vector of coefficients such that $e_{f(x)}^{T}f_{i}=f_{i}(x_{i})$. The second equality holds because $f_{i}=f_{A}$ for all $i$, by definition of the agent-symmetric $\mathbf{f}^{s}$ (76).
@@
    The average over the $n^{2}$ pairs of agents of the scalar products between two variables, each related to any agent :
@@
Then, we prove that, in the symmetrized PEP, any single-agent constraint (Definition 12) can be written with $f_{A}$ and $G_{A}$, independently of $n$. Since the PEP components are all Gram-representable, any of its single-agent constraints for agent $i$ linearly combines local function values $f_{i}(x_{i})$ and scalar products of local variables $x_{i}^{T}y_{i}$, with coefficients that are independent of $n$. Here, $x_{i}$ and $y_{i}$ denote any local vector held by agent $i$. By definition of the symmetric $\mathbf{f}^{s}$ (76), a local function value term can be expressed with $f_{A}$ as
@@
which is well scale-invariant thanks to the $1/n$ factor. Finally, the interpolation constraints for consensus steps with averaging matrices from $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ are given by (65), (66) and (67), from Corollary 4.1. In these constraints, notations hide sums over the agents and can be written with Gram-representable and scale-invariant expressions as
@@
In the end, all the PEP components can well be expressed independently of $n$, with $f_{A}\in\mathbb{R}^{q}$, $G_{A}\in\mathbb{R}^{p\times p}$ and $G_{C}\in\mathbb{R}^{p\times p}$. We refer to Appendix A for explicit expressions of different possible PEP constraints and objectives. Therefore, the SDP PEP can be made smaller and fully independent of the number of agents $n$, by considering these blocks as variables of the problem, instead of the full matrix $G^{s}\in\mathbb{R}^{np\times np}$ and vector $\mathbf{f}^{s}\in\mathbb{R}^{nq}$. ∎
@@ 4.2 Impact of agent symmetry on the worst-case solution
By definition of the fully symmetric PEP solution (76)-(77), in which the Gram matrix has equal diagonal blocks $G_{A}$ and equal off-diagonal blocks $G_{B}$, we can prove symmetry results for the worst-case iterates and functions of the agents. These symmetries open up perspectives with implications for understanding, analyzing and designing distributed algorithms yet to be discovered.
@@
[…]
@@
[…]
@@
By Corollary 6.1, when the agents are all equivalent, there is a worst-case solution of the agent-dependent PEP, which is fully symmetric : $(\mathbf{f},G)=(\mathbf{f}^{s},G^{s})$. When the solution is agent-symmetric, by definition of $G^{s}$ (77), we have
@@
Concerning the local functions, by definition of $\mathbf{f}^{s}$ (76), we have, for all $i,j\in\mathcal{V}$,
[…]
@@ 5 Compact PEP with multiple equivalence classes of agents
While all agents are equivalent (Assumption 1) for many common performance evaluation settings, there are advanced settings for which there are several equivalence classes of agents, see Definition 9. For example, when different groups of agents use different (uncoordinated) step-sizes, function classes, initial conditions, or even different algorithms. It can also happen that the performance measure focuses on a specific group of agents, e.g. the performance of the worst agent. The ability to efficiently and accurately evaluate the performance of distributed optimization algorithms in such advanced settings would enable a more comprehensive analysis and deeper understanding of the algorithms performance.
@@
In this section, we exploit symmetries in different equivalence classes of agents in PEP, revealed in Proposition 6, to write agent-dependent PEPs in a compact form whose size only depends on the number of equivalence classes $r$, and not on the total number of agents $n$. By Proposition 6, we can solve the agent-dependent PEP problem, described in Section 2, restricted to symmetric agent-class solutions $(\mathbf{f}^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$, without impacting the worst-case value. These symmetric solutions $(\mathbf{f}^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$, defined in (74) and (75), depend on the partition $\mathcal{T}$ of the set of agents $\mathcal{V}$ into equivalence classes : $\mathcal{T}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{r}\}$, see Definition 9. A symmetrized solution $(\mathbf{f}^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ has equal blocks for equivalent agents. In this section, we order the agents by equivalence classes in the symmetrized solution $(\mathbf{f}^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ (74)-(75) so that we can partition it into different blocks, associated with the different agent classes. For example, if we have two equivalence classes ($r=2$) : one with two agents and one with one agent, the symmetrized solution can be written as follows :
[…]
where $f_{A}^{1}$, $f_{A}^{2}$, $G_{A}^{1}$, $G_{B}^{1}$, $G_{A}^{2}$, $G_{E}^{12}$ and $G_{E}^{21}$ are blocks defined in Definiton 10. We define new notations to aggregate the blocks of each equivalence class :
$\mathbf{f}^{s}_{\mathcal{T}}$ $=\begin{bmatrix}\mathbf{f}_{\mathcal{V}_{1}}&\mathbf{f}_{\mathcal{V}_{2}}\end{bmatrix}^{T}$ and $\quad G^{s}_{\mathcal{T}}$ $=\begin{bmatrix}G_{\mathcal{V}_{1}}^{s}&G_{\mathcal{V}_{1}\mathcal{V}_{2}}^{s}\\ G_{\mathcal{V}_{2}\mathcal{V}_{1}}^{s}&G_{\mathcal{V}_{2}}^{s}\end{bmatrix}$, (105)
In general, when ordering the agents by equivalence classes, the symmetrized solution $(\mathbf{f}^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ (74)-(75) can be written as
@@
$\mathbf{f}^{s}_{\mathcal{T}}$ $=\begin{bmatrix}\mathbf{f}_{\mathcal{V}_{1}}&\dots&\mathbf{f}_{\mathcal{V}_{r}}\end{bmatrix}^{T}$ $\hskip-5.69046pt\text{where }\mathbf{f}_{\mathcal{V}_{u}}=\mathbf{1}_{n_{u}}\otimes f_{A}^{u}$ (106)
@@
$G^{s}_{\mathcal{T}}$ $=\begin{bmatrix}G_{\mathcal{V}_{1}}^{s}&G_{\mathcal{V}_{1}\mathcal{V}_{2}}^{s}&\dots&G_{\mathcal{V}_{1}\mathcal{V}_{r}}^{s}\\ G_{\mathcal{V}_{2}\mathcal{V}_{1}}^{s}&G_{\mathcal{V}_{2}}^{s}&&\vdots\\ \vdots&&\ddots&\\ G_{\mathcal{V}_{r}\mathcal{V}_{1}}^{s}&\dots&&G_{\mathcal{V}_{r}}^{s}\end{bmatrix}$ where $G_{\mathcal{V}_{u}}^{s}=\begin{cases}G_{A}^{u}&\text{if $n_{u}=1$,}\\ \mathbf{1}_{n_{u}}\mathbf{1}_{n_{u}}^{T}\otimes G_{B}^{u}+(I_{n_{u}}\otimes(G_{A}^{u}-G_{B}^{u}))&\text{if $n_{u}\geq 2$},\end{cases}$ and  $G_{\mathcal{V}_{u}\mathcal{V}_{v}}^{s}=\mathbf{1}_{n_{u}}\mathbf{1}^{T}_{n_{v}}\otimes G_{E}^{uv}$, (107)
@@
and where $f_{A}^{u}\in\mathbb{R}^{q}$, $G_{A}^{u},G_{B}^{u},G_{E}^{uv}\in\mathbb{R}^{p\times p}$ for $u,v=1,\dots,r$ are the blocks repeating in $\mathbf{f}^{s}_{\mathcal{T}}$ and $G^{s}_{\mathcal{T}}$, see Definition 10.
Remark (Limit cases for $r$).
@@
  • $r=n$ : When no agents are equivalent, the partition $\mathcal{T}$ contains $r=n$ equivalence classes of size 1 and then the symmetrized agent-class solution $\mathbf{f}^{s}_{\mathcal{T}}$, $G^{s}_{\mathcal{T}}$, defined in Proposition 6 is identical to the agent-dependent PEP solution $\mathbf{f}$, $G$ :
    $\mathbf{f}^{s}_{\mathcal{T}}=F,\qquad\text{ and }\qquad G^{s}_{\mathcal{T}}=G.$
@@
    Therefore, the resulting PEP formulation is equivalent to the agent-dependent PEP from Section 2 and is not compact as its size scales with $r=n$.
@@
  • $r=1$ : When all the agents are equivalent, the partition $\mathcal{T}$ contains $r=1$ equivalence class of size $n$ and then the symmetrized agent-class solution $\mathbf{f}^{s}_{\mathcal{T}}$, $G^{s}_{\mathcal{T}}$ defined in Proposition 6 is identical to the fully symmetrized solution $\mathbf{f}^{s}$, $G^{s}$ defined in Corollary 6.1
    $\mathbf{f}^{s}_{\mathcal{T}}=\mathbf{f}^{s},\qquad\text{ and }\qquad G^{s}_{\mathcal{T}}=G^{s}.$